NORMAL DISTRIBUTION

A potato chip company creates bags of chips with a mean weight of 144.5 grams and a standard deviation of 7.6 grams. The company will not allow any bag weighing less than 130 grams or more than 160 grams to be sold. What percentage of bags will be sold?
Answer #1

Solution :

=1 - P(130 < x < 160) = P[(130 - 144.5)/ 7.6) < (x - ) /  < (160 - 144.5) / 7.6) ]

= 1 - P(-1.91 < z < 2.04)

= 1 - P(z < 2.04) - P(z < -1.91)

= 1 - P(0.9793 - 0.0281)

= 1 - 0.9512

= 0.0488

Percentage = 4.88%